When 23 grams of methane gas combine with 42 grams of oxygen, 37 grams of water form, plus carbon dioxide. What mass of carbon dioxide is produced?
1
Expert's answer
2020-10-14T05:44:21-0400

CH4(g) + 2 O2 (g) = CO2(g) + 2 H2O (g)


1 : 2 : 1 : 2


23g : 42g : ? : 37g


Therefore : ( 23/42) =0.5 moles


1 : 2


? : 0.5

Therefore : (0.5/2)1 =0.25 moles


If,

37g = 0.5


? = 0.25


( (37)(0.25))/0.5 = 18.5g